OpenStudy (anonymous):

last one of these: one serving of DietMuck is 2/3 of a cup. Jane wants to eat 2/3 of a serving of DietMuck. What fraction of a cup should Jane eat?

OpenStudy (anonymous):

(2/3)x(2/3) = 4/9

OpenStudy (stormfire1):

2/3 * 2/3 = 4/9